OpenStudy (anonymous):

Solve the equation:

OpenStudy (anonymous):

Solve this Equation .... 1/3 - 1/2 = 1

OpenStudy (saifoo.khan):

You missed something?

OpenStudy (anonymous):

pellet, This is the Equation ... T/3 - 1/2 = 1

OpenStudy (saifoo.khan):

No need for pelletting. lol

OpenStudy (anonymous):

I didn't know i couldnt Cuss -___-

OpenStudy (saifoo.khan):

\[\frac{T}{3} = 1 + \frac12\] \[\frac{T}{3} = \frac{1(2) +1}{2}\] \[\frac{T}{3} = \frac{3}{2}\]

OpenStudy (saifoo.khan):

Multiply 3 on both sides, \[T = \frac32 \times 3\] Solve. :)

OpenStudy (anonymous):

Im So confused :( Lmfao

OpenStudy (saifoo.khan):

Where?

OpenStudy (anonymous):

Lol likeeee, Whats the awnser ?

OpenStudy (saifoo.khan):

Direct answer is marked as cheating

OpenStudy (anonymous):

Are you kidding me ....

OpenStudy (saifoo.khan):

Nope read this out, http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I don't have time for reading -____- Im in the middle off a test and stuff,. Lol Ooh weeelll, ill just guess

OpenStudy (anonymous):

the answer could be a) pellet